### v3

@@ 3 Inexact $q$-order regularized proximal Newton method
Combining this with (21) yields that $F(y^{k,j})-F(x^{k})<-\sigma(L_{k,j}/q)\|y^{k,j}\!-x^{k}\|^{q}$, a contradiction to inequality (20). The conclusion then follows.
Remark 3.4.
Fix any $k\!\in\!\mathbb{N}$ with $r(x^{k})\!\neq 0$. By Lemma 3.2 (i), when $L_{k,j}$ is large enough, subproblem (14) has a stationary point $\overline{y}^{k,j}$ with $\Theta_{k,j}(\overline{y}^{k,j})<\Theta_{k,j}(x^{k})$. Next we take a closer look at how to seek an inexact stationary point satisfying (15) with the proximal gradient (PG) method to solve subproblem (14). Fix any $j\in\mathbb{N}$. Its basic iteration is
(22) $y^{l+1}=\mathop{\arg\min}_{y\in\mathbb{X}}\,\langle\nabla\!f_{k,j}(y^{l}),y-y^{l}\rangle+\frac{1}{2\tau_{l}}\|y-y^{l}\|^{2}+g(y)\ \ {\rm with}\ \ y^{0}=x^{k},$
where $\tau_{l}>0$ is the step-size depending on the Lipschitz constant of $\nabla\!f_{k,j}$. We assume that $\underline{\tau}\leq\tau_{l}$ with some $\underline{\tau}>0$ for all $l$. By the expression of $f_{k,j}$, for any $y,y^{\prime}\in\mathbb{X}$,
$\|\nabla\!f_{k,j}(y^{\prime})-\nabla\!f_{k,j}(y)\|$ $\leq\|\nabla^{2}\!f(x^{k})\|\|y^{\prime}-y\|+L_{k,j}\|y^{\prime}-x^{k}\|^{q-2}\|y^{\prime}-y\|$
(23) $\quad+L_{k,j}\big{|}\|y^{\prime}-x^{k}\|^{q-2}-\|y-x^{k}\|^{q-2}\big{|}\|y-x^{k}\|.$
From [45, Lemma 3] and the optimal conditions of (22), for each $l\in\mathbb{N}$, it holds that
$r_{k,j}(y^{l+1})$ $\leq\frac{1}{\tau_{l}}\|y^{l+1}\!-\!{\rm prox}_{\tau_{l}g}(y^{l+1}\!-\!\tau_{l}\nabla\!f_{k,j}(y^{l+1}))\|$
$=\frac{1}{\tau_{l}}\|{\rm prox}_{\tau_{l}g}(y^{l}\!-\!\tau_{l}\nabla\!f_{k,j}(y^{l}))-{\rm prox}_{\tau_{l}g}(y^{l+1}\!-\!\tau_{l}\nabla\!f_{k,j}(y^{l+1}))\|$
$\leq\frac{1}{\tau_{l}}\|y^{l}-y^{l+1}\|+\|\nabla\!f_{k,j}(y^{l})\!-\!\nabla\!f_{k,j}(y^{l+1})\|$
$\leq(\tau_{l}^{-1}\!+\!\|\nabla^{2}\!f(x^{k})\|)\|y^{l+1}\!-\!y^{l}\|+L_{k,j}\|y^{l+1}\!-\!x^{k}\|^{q-2}\|y^{l+1}\!-\!y^{l}\|$
(24) $\quad+L_{k,j}|\|y^{l+1}\!-\!x^{k}\|^{q-2}-\|y^{l}\!-\!x^{k}\|^{q-2}|\|y^{l}-x^{k}\|$
where the last inequality is due to (3.4). From the convergence analysis of the line-search PG method [26], the objective value sequence $\{\Theta_{k,j}(y^{l})\}_{l\in\mathbb{N}}$ is sufficiently descent and $\|y^{l+1}\!-\!y^{l}\|\to 0$ as $l\to\infty$, and under the KL property of the objective functions of subproblems, the sequence $\{y^{l}\}_{l\in\mathbb{N}}$ is convergent with limit, say, $y^{*}$. Assume that $y^{*}\neq x^{k}$ (if not, from the first-order optimality condition of (22), we conclude that $x^{k}$ is a stationary point of (1)). By the continuous differentiability of the function $z\mapsto\|z-x^{k}\|$ at $y^{l}$ for all sufficiently large $l$, one can verify that the term on the right hand side of (3.4) approaches to 00 as $l\to\infty$, so it is easily controlled from above by $\varrho L_{k,j}\|y^{l+1}\!-\!x^{k}\|^{q-1}$ for all sufficiently large $l$. Thus, the line-search PG method with the stopping condition $\frac{1}{\tau_{l}}\|y^{l}\!-\!y^{l+1}\|+\|\nabla\!f_{k,j}(y^{l})\!-\!\nabla\!f_{k,j}(y^{l+1})\|\leq\varrho L_{k,j}\|y^{l+1}\!-\!x^{k}\|^{q-1}$ will return $y^{l+1}$ satisfying condition (15). The prerequisite for the above analysis is that the line-search PG method works. When the subproblems have a bad scaling, i.e., $\tau_{l}^{-1}\!+\!\|\nabla^{2}\!f(x^{k})\|$ is larger, the line-search PG method will have a slow convergence. In this case, a second-order algorithm is suggested as the inner solver.
For convenience, in the rest of this paper, we let $\{x^{k}\}_{k\in\mathbb{N}}$ and $\{y^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by Algorithm 1 with $\epsilon=0$. Next we take a closer look at the properties of the iterate and objective value sequences of Algorithm 1.
@@
where the second inequality is by Proposition 3.5 (i), and the third one is obtained by using (19) with $x=y^{k,j_{k}\!-\!1}$. From Assumption 2 (i), the continuity of $\nabla^{2}\!f$ on ${\rm dom}\,g$ and $\lim_{\mathcal{K}\ni k\to\infty}\widetilde{L}_{k}=\infty$, the above inequality implies that $\lim_{\mathcal{K}\ni k\to\infty}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|=0$. Then, for each $k\in\mathcal{K}$ large enough, using the mean-value theorem and Assumption 2 (ii) and following the same arguments as those for [40, Lemma 4.2 (iv)] leads to $(1\!-\!\sigma)(\widetilde{L}_{k}/q)\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|^{q}\leq\frac{c_{0}}{2}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|^{3}$ for some $c_{0}>0$, which is impossible because $\lim_{\mathcal{K}\ni k\to\infty}\widetilde{L}_{k}=\infty$, $y^{k,j_{k}\!-\!1}\neq x^{k}$ by (25) and $\lim_{\mathcal{K}\ni k\to\infty}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|=0$.
@@ 4.1 Global convergence
By [42, Exercise 10.7], $\partial\Phi(w)=\{(\xi;-\xi)\ |\ \xi\in\partial F(y-v)\}$ for any $w=(y,v)$. Hence, if $w^{*}=(y^{*},v^{*})$ is a critical point of $\Phi$, then $x^{*}=y^{*}-v^{*}$ is a critical point of $F$; and conversely, if $x^{*}$ is a critical point of $F$, then $w^{*}=(x^{*},0)$ is a critical point of $\Phi$. By the expression of $\Phi$ and the proof of [29, Theorem 3.2], the following fact holds.
@@ 4.2 Local convergence rates
By Fact 1, if $F$ is a KL function of exponent $\theta\in[0,1)$, then $\Phi$ is a KL function of exponent $\theta\in[0,1)$. Based on the proof of Theorem 4.3, by using Proposition 3.5 and following the similar arguments to those of [40, Theorem 3.2], one can achieve the following local convergence rate result.
@@
Next we use the relation in (40) to establish the superlinear rate of the residual sequence $\{r(x^{k})\}_{k\in\mathbb{N}}$ and the iterate sequence $\{x^{k}\}_{k\in\mathbb{N}}$. Such a relation was used in the local convergence rate analysis of inexact proximal Newton-type methods [48, 34, 31]. It is worth emphasizing that our proof is not a direct extension of theirs because the direction $y^{k}-x^{k}$ in the aforementioned works is obtained by minimizing a strongly convex function. When $g\equiv 0$ and the subproblems are solved exactly, such a result was achieved in [49, Lemma 1], but its analysis is inapplicable to our algorithm.
@@
From $\overline{x}\in\omega(x^{0})$, there exists $\mathcal{K}\subset\mathbb{N}$ such that $\lim_{\mathcal{K}\ni k\to\infty}x^{k}=\overline{x}$. Along with the given assumption, ${\rm dist}(x^{k},\mathcal{X}^{*})\leq\kappa[r(x^{k})]^{\gamma}$ for all $k\in\mathcal{K}$ large enough. Passing the limit $\mathcal{K}\ni k\to\infty$ and using Proposition 3.6 (ii) leads to $\overline{x}\in\mathcal{X}^{*}$. For each $k\geq\widetilde{k}$ where $\widetilde{k}$ is the same as in Proposition 4.8, since $x^{k}\!\in\Gamma$ and $\overline{x}\in\omega(x^{0})\!\subset\Gamma$, invoking the nonexpansiveness of ${\rm prox}_{g}$ results in
[…]
@@
(40)${{\leq}}\eta(1\!+\!\varsigma)\sum_{l=\overline{k}_{2}}^{k}{\rm dist}(x^{l},\mathcal{X}^{*})\leq\eta\kappa(1\!+\!\varsigma)\sum_{l=\overline{k}_{2}}^{k}[r(x^{l})]^{\gamma}$
(48)${{\leq}}\eta\kappa(1\!+\!\varsigma)[r(x^{\overline{k}_{2}})]^{\gamma}\sum_{l=\overline{k}_{2}}^{k}\widetilde{\sigma}^{l-\overline{k}_{2}}\leq\frac{\eta\kappa(1\!+\!\varsigma)}{1-\widetilde{\sigma}}[r(x^{\overline{k}_{2}})]^{\gamma}$
@@
Theorem 4.10 states that the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to a second-order stationary point with $Q$-superlinear rate of order $\gamma(q\!-\!1)$, which is specified as $2$ for $q=3$ and $\gamma\!=1$, under the strict continuity of $\nabla^{2}\!f$ on an open convex set $\mathcal{N}$ with $\mathcal{O}\supset\mathcal{N}\supset\omega(x^{0})$ and the local Hölderian error bound of order $\gamma\in(\frac{1}{q-1},1]$ on $\mathcal{X}^{*}$. To the best of our knowledge, even for nonsmooth convex composite optimization, this is the first quadratic convergence rate of the iterate sequence for CR methods under a local Lipschitzian error bound, which is equivalent to the KL property of $F$ with exponent $1/2$ by [31, Proposition 2 (i)] for convex composite optimization.
@@
Fix any $q\in(2,3]$ and $\overline{x}\in\omega(x^{0})$. Suppose that $F$ is a KL function, that the assumptions of Theorem 4.10 hold at $\overline{x}$ with $\gamma\in(\frac{1}{q-1},1]$, and that there exist $\widetilde{\varepsilon}\in(0,\widehat{\varepsilon})$ and $\widehat{\kappa}\in(0,\kappa^{-\frac{1}{\gamma}})$ such that for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$,
(52) $F(x)\geq F(\overline{x})-\frac{\gamma\widehat{\kappa}}{1+\gamma}[{\rm dist}(x,\mathcal{S}^{*})]^{\frac{1+\gamma}{\gamma}}.$
Then, for all sufficiently large $k$, the following inequality holds
@@
From [31, Remark 1 (b) [^3] The conclusion of [31, Remark 1(b)] still holds under the subdifferential continuity of $F$, which is implied by Assumption 1.], $F(z^{*})\leq F(\overline{x})=\overline{F}$ for all $z^{*}\in\mathcal{S}\cap\mathbb{B}(\overline{x},\widetilde{\varepsilon})$ (if necessary by shrinking $\widetilde{\varepsilon}$). As $\mathcal{X}^{*}\subset\mathcal{S}^{*}$, the local error bound condition implies that ${\rm dist}(x,\mathcal{S}^{*})\leq\kappa[r(x)]^{\gamma}$ for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$, which by [31, Lemma 1] is equivalent to the $\gamma$-subregularity of $\partial F$ at $(\overline{x},0)$ with modulus $\kappa$. Together with (52) and [33, Theorem 3.4], there exists $\beta_{0}>0$ such that for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$,
@@
(15)${{\leq}}\varrho L_{k}\|y^{k}-x^{k}\|^{q-1}+{\rm dist}(x^{k+1},\mathcal{X}^{*})$ )
(50)${{\leq}}\varrho L_{k}\|y^{k}-x^{k}\|^{q-1}+\kappa\widehat{c}^{\gamma}{\rm dist}(x^{k},\mathcal{X}^{*})^{\gamma(q-1)}$
@@
where the last inequality is due to Proposition 3.6 (i). As $\lim_{k\to\infty}{\rm dist}(x^{k+1},\mathcal{X}^{*})=0$, we have $x^{k+1,*}\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$ and hence $F(x^{k+1,*})\leq\overline{F}$ for all sufficiently large $k$. Together with equation (32), we have $\Phi(w^{k})\!-\!\overline{F}\leq F(y^{k}\!-\!v^{k})-F(x^{k+1,*})$ for all sufficiently large $k$. Invoking the second inequality of Lemma 4.6 with $x^{*}=x^{k+1,*}$ and the above (4.14), for all sufficiently large $k$, it holds that
@@
(a) Theorem 4.13 shows that the objective value sequence has the same local convergence rate as the iterate sequence, but it requires the local Hölderian error bound to hold at $\overline{x}\in\omega(x^{0})$ for which $F$ satisfies the lower second-order growth in (52). This condition automatically holds if $\overline{x}$ is locally optimal.
@@ 4.3 Complexity bounds
Suppose that Assumptions 1-2 (i) holds, and that $\nabla^{2}\!f$ is strictly continuous on a compact convex set $\Omega$ containing $\{x^{k}\}_{k\in\mathbb{N}}\cup\{y^{k}\}_{k\in\mathbb{N}}$. Then, $r(x^{k+1})\leq\epsilon$ for all $k\geq K\!:=\!\lfloor q(\sigma L_{m})^{-1}\min\{\widehat{\eta}^{\frac{q}{2}},\widehat{\eta}^{\frac{q}{q-1}}\}(F(x^{0})\!-\!\overline{F})\epsilon^{-\frac{q}{q-1}}\rfloor$ with $\widehat{\eta}\!=\!(\varrho(3\!+\!L_{g})\!+\!1)\widehat{L}\!+\!{L_{H}}/{2}$, where $L_{H}$ denotes the Lipschitz constant of $\nabla^{2}\!f$ on the set $\Omega$.
@@
Since $\nabla^{2}\!f(\cdot)$ is assumed to be strictly continuous on the compact convex set $\Omega$, it is Lipschitz continuous on $\Omega$ by [42, Theorem 9.2]. We first derive an upper bound of $r(x^{k+1})$ in terms of $\|y^{k}\!-\!x^{k}\|$. It is easy to check (4.11) holds for all $k\in\mathbb{N}$. From the Lipschitz continuity of $\nabla^{2}\!f(\cdot)$ on $\Omega$ and $\{(x^{k},y^{k})\}_{k\in\mathbb{N}}\subset\Omega$, inequality (33) holds for all $k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$, it holds that
@@
By combining this inequality with (4.11), for each $k\in\mathbb{N}$, it holds that
@@
where the second inequality comes from $\|v^{k}\|\leq\varrho L_{k}\|y^{k}-x^{k}\|^{q-1}$ in (15) and $L_{k}\leq\widehat{L}$ in Proposition 3.6 (i). As $r(x^{K+1})\leq\epsilon$ and Algorithm 1 stops at the $(K\!+\!1)$th iteration, we have $r(x^{i})>\epsilon$ for $i=1,\ldots,K$. Fix any $i\in\{1,\ldots,K\}$. When $\|y^{i-1}\!-\!x^{i-1}\|>1$, combining the above (4.17) with $r(x^{i})>\epsilon$ and the definition of $\widehat{\eta}$ leads to $\|y^{i-1}\!-\!x^{i-1}\|\geq\widehat{\eta}^{-1/2}\epsilon^{1/2}\geq\widehat{\eta}^{-1/2}\epsilon^{1/(q-1)}$. When $\|y^{i-1}\!-\!x^{i-1}\|\leq 1$, the above (4.17) along with $r(x^{i})>\epsilon$ and the definition of $\widehat{\eta}$ implies that $\|y^{i-1}\!-\!x^{i-1}\|\geq\widehat{\eta}^{-1/(q-1)}\epsilon^{1/(q-1)}$. Thus, $\|y^{i-1}\!-\!x^{i-1}\|\geq\min\{\widehat{\eta}^{-1/2},\widehat{\eta}^{-1/(q-1)}\}\epsilon^{1/(q-1)}$. Together with Proposition 3.5 (i), it follows that
$F(x^{i-1})-F(x^{i})\geq\frac{\sigma L_{m}}{q}\|y^{i-1}\!-\!x^{i-1}\|^{q}\geq\frac{\sigma L_{m}}{q}\min\{\widehat{\eta}^{-q/2},\widehat{\eta}^{-q/(q-1)}\}\epsilon^{\frac{q}{q-1}}.$
@@
$F(x^{0})-\overline{F}\geq\sum_{i=1}^{K}(F(x^{i-1})\!-\!F(x^{i}))\geq K\frac{\sigma L_{m}}{q}\min\{\widehat{\eta}^{-q/2},\widehat{\eta}^{-q/(q-1)}\}\epsilon^{\frac{q}{q-1}},$

### v4

@@ 3 Inexact $q$-order regularized proximal Newton method
Combining this with (17) yields that $F(y^{k,j})-F(x^{k})\leq-\sigma(L_{k,j}/q)\|y^{k,j}\!-x^{k}\|^{q}$, a contradiction to inequality (16). The conclusion then follows.
For convenience, in the rest of this paper, we let $\{x^{k}\}_{k\in\mathbb{N}}$ and $\{y^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by Algorithm 1. Next we take a closer look at the properties of the iterate and objective value sequences of Algorithm 1.
@@
where the second inequality is by Proposition 3.4 (i), and the third one is obtained by using (15) with $x=y^{k,j_{k}\!-\!1}$. From Assumption 2 (i), the continuity of $\nabla^{2}\!f$ on ${\rm dom}\,g$ and $\lim_{\mathcal{K}\ni k\to\infty}\widetilde{L}_{k}=\infty$, the above inequality implies that $\lim_{\mathcal{K}\ni k\to\infty}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|=0$. Then, for each $k\in\mathcal{K}$ large enough, using the mean-value theorem and Assumption 2 (ii) and following the same arguments as those for [39, Lemma 4.2 (iv)] leads to $(1\!-\!\sigma)(\widetilde{L}_{k}/q)\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|^{q}\leq\frac{c_{0}}{2}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|^{3}$ for some $c_{0}>0$, which is impossible because $\lim_{\mathcal{K}\ni k\to\infty}\widetilde{L}_{k}=\infty$, $y^{k,j_{k}\!-\!1}\neq x^{k}$ by (18) and $\lim_{\mathcal{K}\ni k\to\infty}\|y^{k,j_{k}\!-\!1}\!-\!x^{k}\|=0$.
@@ 4.1 Global convergence
By [41, Exercise 10.7], $\partial\Phi(w)=\{(\xi;-\xi)\ |\ \xi\in\partial F(y-v)\}$ for any $w=(y,v)$. Hence, if $w^{*}=(y^{*},v^{*})$ is a critical point of $\Phi$, then $x^{*}=y^{*}-v^{*}$ is a critical point of $F$; and conversely, if $x^{*}$ is a critical point of $F$, then $w^{*}=(x^{*},0)$ is a critical point of $\Phi$. By the expression of $\Phi$ and the proof of [28, Theorem 3.2], the following fact holds.
@@ 4.2 Local convergence rates
By Fact 1, if $F$ is a KL function of exponent $\theta\in[0,1)$, then $\Phi$ is a KL function of exponent $\theta\in[0,1)$. Based on the proof of Theorem 4.3, by using Proposition 3.4 and following the similar arguments to those of [39, Theorem 3.2], one can achieve the following local convergence rate result.
@@
Next we use the relation in (33) to establish the superlinear rate of the residual sequence $\{r(x^{k})\}_{k\in\mathbb{N}}$ and the iterate sequence $\{x^{k}\}_{k\in\mathbb{N}}$. Such a relation was used in the local convergence rate analysis of inexact proximal Newton-type methods [46, 33, 30]. It is worth emphasizing that our proof is not a direct extension of theirs because the direction $y^{k}-x^{k}$ in the aforementioned works is obtained by minimizing a strongly convex function. When $g\equiv 0$ and the subproblems are solved exactly, such a result was achieved in [47, Lemma 1], but its analysis is inapplicable to our algorithm.
@@
From $\overline{x}\in\omega(x^{0})$, there exists $\mathcal{K}\subset\mathbb{N}$ such that $\lim_{\mathcal{K}\ni k\to\infty}x^{k}=\overline{x}$. Along with the given assumption, ${\rm dist}(x^{k},\mathcal{X}^{*})\leq\kappa[r(x^{k})]^{\gamma}$ for all $k\in\mathcal{K}$ large enough. Passing the limit $\mathcal{K}\ni k\to\infty$ and using Proposition 3.5 (ii) leads to $\overline{x}\in\mathcal{X}^{*}$. For each $k\geq\widetilde{k}$ where $\widetilde{k}$ is the same as in Proposition 4.8, since $x^{k}\!\in\Gamma$ and $\overline{x}\in\omega(x^{0})\!\subset\Gamma$, invoking the nonexpansiveness of ${\rm prox}_{g}$ and equation (22) results in
(37) $r(x^{k})=|r(x^{k})-r(\overline{x})|\leq 2\|x^{k}-\overline{x}\|+\|\nabla\!f(x^{k})-\nabla\!f(\overline{x})\|\leq(2\!+\!L_{g})\|x^{k}\!-\!\overline{x}\|.$
@@
(33)${{\leq}}\eta(1\!+\!\varsigma){\textstyle\sum_{l=\overline{k}_{2}}^{k}}{\rm dist}(x^{l},\mathcal{X}^{*})\leq\eta\kappa(1\!+\!\varsigma){\textstyle\sum_{l=\overline{k}_{2}}^{k}}[r(x^{l})]^{\gamma}$
(41)${{\leq}}\eta\kappa(1\!+\!\varsigma)[r(x^{\overline{k}_{2}})]^{\gamma}{\textstyle\sum_{l=\overline{k}_{2}}^{k}}\widetilde{\sigma}^{l-\overline{k}_{2}}\leq\frac{\eta\kappa(1\!+\!\varsigma)}{1-\widetilde{\sigma}}[r(x^{\overline{k}_{2}})]^{\gamma}$
@@
Theorem 4.10 states that the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ converges to a second-order stationary point with $Q$-superlinear rate of order $\gamma(q\!-\!1)$, which is specified as $2$ for $q=3$ and $\gamma\!=1$, under the strict continuity of $\nabla^{2}\!f$ on an open convex set $\mathcal{N}$ with $\mathcal{O}\supset\mathcal{N}\supset\omega(x^{0})$ and the local Hölderian error bound of order $\gamma\in(\frac{1}{q-1},1]$ on $\mathcal{X}^{*}$. To the best of our knowledge, even for nonsmooth convex composite optimization, this is the first quadratic convergence rate of the iterate sequence for CR methods under a local Lipschitzian error bound, which is equivalent to the KL property of $F$ with exponent $1/2$ by [30, Proposition 2 (i)] for convex composite optimization.
@@
Fix any $q\in(2,3]$ and $\overline{x}\in\omega(x^{0})$. Suppose that $F$ is a KL function, that the assumptions of Theorem 4.10 hold at $\overline{x}$ with $\gamma\in(\frac{1}{q-1},1]$, and that there exist $\widetilde{\varepsilon}\in(0,\widehat{\varepsilon})$ and $\widehat{\kappa}\in(0,\kappa^{-\frac{1}{\gamma}})$ such that $F(x)\geq F(\overline{x})-\frac{\gamma\widehat{\kappa}}{1+\gamma}[{\rm dist}(x,\mathcal{S}^{*})]^{\frac{1+\gamma}{\gamma}}$ for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$. Then, for all sufficiently large $k$, the following inequality holds
@@
From [30, Remark 1 (b)], $F(z^{*})\leq F(\overline{x})=\overline{F}$ for all $z^{*}\in\mathcal{S}\cap\mathbb{B}(\overline{x},\widetilde{\varepsilon})$ (if necessary by shrinking $\widetilde{\varepsilon}$). As $\mathcal{X}^{*}\subset\mathcal{S}^{*}$, the local error bound condition implies that ${\rm dist}(x,\mathcal{S}^{*})\leq\kappa[r(x)]^{\gamma}$ for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$, which by [30, Lemma 1] is equivalent to the $\gamma$-subregularity of $\partial F$ at $(\overline{x},0)$ with modulus $\kappa$. Together with the given assumption on $F$ and [32, Theorem 3.4], there exists $\beta_{0}>0$ such that for all $x\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$,
@@
[…]
@@
where the last inequality is due to Proposition 3.5 (i). As $\lim_{k\to\infty}{\rm dist}(x^{k+1},\mathcal{X}^{*})=0$, we have $x^{k+1,*}\in\mathbb{B}(\overline{x},\widetilde{\varepsilon})$, and then $F(x^{k+1,*})\leq\overline{F}$ for all $k$ large enough. Together with (25), for all $k$ large enough, $\Phi(w^{k})\!-\!\overline{F}\leq F(y^{k}\!-\!v^{k})-F(x^{k+1,*})$. Invoking the second inequality of Lemma 4.6 with $x^{*}=x^{k+1,*}$ and (4.14), for all $k$ large enough,
@@
(a) Theorem 4.13 shows that the objective value sequence has the same local convergence rate as the iterate sequence, but it requires the local Hölderian error bound to hold at $\overline{x}\in\omega(x^{0})$ for which $F$ satisfies the lower second-order growth. This condition automatically holds when $\overline{x}$ is locally optimal.
@@ 4.3 Complexity bounds
Suppose that Assumptions 1-2 (i) holds, and that $\nabla^{2}\!f$ is strictly continuous on a compact convex set $\Omega$ containing $\{x^{k}\}_{k\in\mathbb{N}}\cup\{y^{k}\}_{k\in\mathbb{N}}$. Then, for any given $\epsilon\in(0,1)$, $r(x^{K+1})\leq\epsilon$ with $K\!:=\!\lfloor q(\sigma L_{m})^{-1}\min\{\widehat{\eta}^{\frac{q}{2}},\widehat{\eta}^{\frac{q}{q-1}}\}(F(x^{0})\!-\!\overline{F})\epsilon^{-\frac{q}{q-1}}\rfloor$ for $\widehat{\eta}\!=\!(\varrho(3\!+\!L_{g})\!+\!1)\widehat{L}\!+\!{L_{H}}/{2}$, where $L_{H}$ is the Lipschitz constant of $\nabla^{2}\!f$ on $\Omega$.
@@
Since $\nabla^{2}\!f(\cdot)$ is assumed to be strictly continuous on the compact convex set $\Omega$, it is Lipschitz continuous on $\Omega$ by [41, Theorem 9.2]. We first derive an upper bound of $r(x^{k+1})$ in terms of $\|y^{k}\!-\!x^{k}\|$. It is easy to check (4.11) holds for all $k\in\mathbb{N}$. From the Lipschitz continuity of $\nabla^{2}\!f(\cdot)$ on $\Omega$ and $\{(x^{k},y^{k})\}_{k\in\mathbb{N}}\subset\Omega$, inequality (26) holds for all $k\in\mathbb{N}$. Then, for all $k\in\mathbb{N}$, it holds that
@@
By combining this inequality with (4.11) and (11), for each $k\in\mathbb{N}$, it holds that
@@
where the second inequality is due to $L_{k}\leq\widehat{L}$ by Proposition 3.5 (i). As $r(x^{K+1})\leq\epsilon$ and Algorithm 1 stops at the $(K\!+\!1)$th iteration, we have $r(x^{i})>\epsilon$ for $i=1,\ldots,K$. Fix any $i\in\{1,\ldots,K\}$. When $\|y^{i-1}\!-\!x^{i-1}\|>1$, combining the above (4.17) with $r(x^{i})>\epsilon$ and the definition of $\widehat{\eta}$ leads to $\|y^{i-1}\!-\!x^{i-1}\|\geq\widehat{\eta}^{-1/2}\epsilon^{1/2}\geq\widehat{\eta}^{-1/2}\epsilon^{1/(q-1)}$. When $\|y^{i-1}\!-\!x^{i-1}\|\leq 1$, the above (4.17) along with $r(x^{i})>\epsilon$ and the definition of $\widehat{\eta}$ implies that $\|y^{i-1}\!-\!x^{i-1}\|\geq\widehat{\eta}^{-1/(q-1)}\epsilon^{1/(q-1)}$. Thus, $\|y^{i-1}\!-\!x^{i-1}\|\geq\min\{\widehat{\eta}^{-1/2},\widehat{\eta}^{-1/(q-1)}\}\epsilon^{1/(q-1)}$. Together with Proposition 3.4 (i), it follows that
$F(x^{i-1})-F(x^{i})\geq\sigma L_{m}q^{-1}\|y^{i-1}\!-\!x^{i-1}\|^{q}\geq\sigma L_{m}q^{-1}\min\{\widehat{\eta}^{-q/2},\widehat{\eta}^{-q/(q-1)}\}\epsilon^{\frac{q}{q-1}}.$
@@
$F(x^{0})-\overline{F}\geq{\textstyle\sum_{i=1}^{K}}(F(x^{i-1})\!-\!F(x^{i}))\geq K\sigma L_{m}q^{-1}\min\{\widehat{\eta}^{-q/2},\widehat{\eta}^{-q/(q-1)}\}\epsilon^{\frac{q}{q-1}},$